Unit 30 Homework Assignment  Physics141-Summer 2009

Unit 30 homework after session 1

• Read chapter 35 in the text book Understanding Physics.

• Work textbook problems 35-9, 35-13, 35-35. Unit 30 homework after session 2

• Read chapter 35 in the text book Understanding Physics..

• Work textbook problems 35-46, 35-51, and 35-52.

All Unit 30 homework problems and the unit 30 activity guide are due on the first day of unit 31.